### v2

# Constraining electromagnetic couplings of ultralight scalars from compact stars

Abstract
If an ultralight scalar interacts with the electromagnetic fields of a compact rotating star, then a long-range scalar field is developed outside the star. The Coulomb-like profile of the scalar field to the leading order is equivalent to an effective scalar charge on the star. In a binary star system, the scalar-induced charge would result in a long-range force between the stars, with the scalar field acting as the mediator. The scalar-photon interactions would modify Maxwell’s equations for electromagnetic fields in vacuum, resulting in a modified dispersion relation. This could be observed as an apparent redshift for photons emitted by such sources. The scalar field would also induce additional electric and magnetic fields and hence affect the electromagnetic energy radiated from such compact objects. A scalar field sourced by time-varying electromagnetic fields can also carry away energy from a compact star in the form of radiation, and hence contribute to its spin-down luminosity. We constrain the scalar-photon coupling from the measurements of the electromagnetic radiation of a compact star and from its spin-down luminosity, using the Crab pulsar, the soft gamma repeater SGR 1806-20, and the gamma ray burst GRB 080905A. We also project the prospective bounds on the coupling from future measurements of the long-range force between two compact stars in a binary such as PSR J0737-3039, and from the apparent redshifts of compact stars. Future advances in precision-clock sensitivity and targeted observations of stars with strong surface magnetic fields, large radii, and low-frequency emission can substantially tighten these coupling limits.
preprint: TIFR-TH/25-1
I Introduction
Neutron stars (NSs) — or pulsars — act as remarkable cosmic laboratories for exploring the mysteries of the Universe. They play a crucial role in generating gravitational waves (GWs), as evidenced by the GW170817 event LIGOScientific:2017vwq that has paved the way for advancements in multi-messenger astronomy Cowperthwaite:2017dyu . These dense, rotating, magnetized objects emit radio waves so regularly that they behave like cosmic clocks. The typical mass of a NS is $1.4~M_{\odot}$ and its radius is $10-20~\mathrm{km}$. The magnetic field of the NS is dipolar and its strength is about $10^{12}~\mathrm{G}$ Gold:1968zf ; Goldreich:1969sb ; Sturrock:1971zc ; Ruderman:1975ju ; Weekes:1989tc ; Lattimer:2004pg . If the magnetic field is even stronger $(\gtrsim 10^{15}~\mathrm{G})$, then the compact object is called a magnetar Duncan:1992hi ; Usov:1992zd ; Stella:2005yz ; Turolla:2015mwa ; Kaspi:2017fwg ; Beniamini:2017ilu ; Esposito:2018gvp ; CHIMEFRB:2020abu ; Lin:2020vlo ; DallOsso:2021xbv .
Compact stars (NSs and magnetars) also serve as probes to search for the dark matter (DM) in the Universe Raffelt:1999tx ; Raffelt:2006cw ; Bramante:2023djs ; Caputo:2024oqc . Results from the Planck satellite suggest that the energy density of DM is about five times that of the visible matter Planck:2018nkj . The weakly-interacting massive particle (WIMP) motivated by the theory of supersymmetry has been one of the leading candidates for DM Jungman:1995df . However, constraints on WIMPs from direct detection experiments Roszkowski:2017nbc ; XENON:2018voc ; XENON:2020gfr ; LZ:2022lsv and the small scale structure of the Universe Bullock:2017xww motivate us to study alternative candidates for DM. Ultralight DM is one such promising candidate, where sub-eV mass range particles can account for the present DM relic density of the universe, at the same time staying consistent with the direct search experiments and cosmological observations Hu:2000ke ; Marsh:2015wka ; Hui:2016ltb ; Robles:2018fur ; Amin:2022pzv . If such a DM candidate has mass as low as $10^{-22}~\mathrm{eV}$, its de Broglie wavelength would be of the order of the size of a dwarf galaxy $(1-2~\mathrm{kpc})$. The number density of ultralight DM within this de Broglie wavelength is $10^{30}/\mathrm{cm^{3}}$ for the local DM density $\rho_{\odot}\sim 0.4~\mathrm{GeV}/\mathrm{cm^{3}}$. The presence of such a large number density implies that DM oscillates coherently in a wave-like manner or exhibits long-range behavior, potentially forming a Bose-Einstein condensate Boehmer:2007um ; Li:2013nal ; KumarPoddar:2019jxe ; DiGiovanni:2022mkn ; Davoudiasl:2024grq ; Poddar:2024thb ; Dutta:2024vzw . The ultralight DM can be scalar Khmelnitsky:2013lxt ; Hui:2016ltb ; Badurina:2021lwr ; Tretiak:2022ndx ; Matos:2023usa , pseudoscalar Zhang:2016uiy ; Hook:2017psm ; Hook:2018iia ; Dror:2020zru ; Chadha-Day:2021szb ; Winch:2024mrt ; Li:2024bbe , vector Dror:2018pdh ; PPTA:2022eul ; Fedderke:2022ptm ; Chase:2023puj ; Chowdhury:2023xvy , or tensor Brito:2020lup ; Wu:2023dnp ; Guo:2023gfc ; some such particles are also motivated from string/M theory Witten:1984dg ; Battye:1994au ; Yamaguchi:1998gx ; Svrcek:2006yi ; Arvanitaki:2009fg .
In addition to its gravitational interactions, if the DM interacts with the Standard Model (SM) particles with very small interaction strengths (allowed by the current data), then precision measurements at the existing and forthcoming experiments can either detect or constrain its properties. No observations or experiments have found the nature of DM so far. However, there are several tests which put constraints on ultralight DM, for example, gravity tests Wagner:2012ui ; KumarPoddar:2019jxe ; KumarPoddar:2019ceq ; KumarPoddar:2020kdz ; Poddar:2020qft ; Poddar:2021sbc ; Poddar:2021ose ; Poddar:2023pfj , magnetometer searches Budker:2013hfa ; Kim:2021eye , Lyman-$\alpha$ observations Nori:2018pka ; Rogers:2020ltq , search for black hole (BH) superradiance Arvanitaki:2010sy ; Baryakhtar:2017ngi ; Brito:2017zvb , variation of fundamental constants Flambaum:2007my ; Stadnik:2014tta ; Leefer:2016xfu ; Kaplan:2022lmz ; Oswald:2021vtc ; Brzeminski:2022sde ; Sherrill:2023zah ; Bloch:2023uis , cosmic microwave background (CMB) observations Hlozek:2014lca ; Hlozek:2017zzf , and more Joshipura:2003jh ; Bandyopadhyay:2006uh ; Huang:2018cwo ; Dror:2020fbh ; Alonso-Alvarez:2021pgy ; Gherghetta:2023myo ; Alonso-Alvarez:2023tii ; Arvanitaki:2014wva ; Arvanitaki:2016qwi ; Kopp:2018jom ; Tsai:2021irw ; Dentler:2021zij ; Dalal:2022rmp . The existing bounds on the coupling of ultralight DM with photons, as determined from different experiments, are summarized in AxionLimits .
The phenomenology of ultralight scalar, pseudoscalar, vector, and tensor fields is remarkably diverse, and numerous studies over the years have explored their potential signatures in cosmic laboratories. The dilaton and axion field profiles in different string gravity models of BH have been discussed in Campbell:1990ai ; Campbell:1991kz ; Campbell:1991rz ; Mignemi:1992pm , where the field is sourced by the Chern-Simon and Gauss-Bonnet terms. In this paper, we consider the scenario where an ultralight scalar $\phi$ (which need not be the DM) interacts with the CP-even electromagnetic (EM) current $F_{\mu\nu}F^{\mu\nu}$. As we shall show later, such an interaction, in the presence of the large EM fields near the surface of a rotating compact star, leads to a long-range scalar field, [^3] Note that ultralight pseudoscalars such as axions may interact with the charge-parity (CP)-odd EM current $F_{\mu\nu}\tilde{F}^{\mu\nu}$. However, the resultant pseudoscalar field goes as $a\sim\cos\theta/r^{2}$ Mohanty:1993nh , where $\theta$ is the polar angle for a rotating magnetized compact star, i.e., it falls faster than the scalar field as one moves away from the source. The influence of an axion background on EM radiation is examined in Harari:1992ea ; Krasnikov:1996bm ; McDonald:2019wou ; Sokolov:2022fvs ; Domcke:2023bat . $\phi\sim 1/r$. This scalar field, in turn, induces additional electric and magnetic fields around the source. We explore four kinds of effects of such a scalar field:
  • The scalar interaction with the EM field of a compact star alters Maxwell’s equations Domcke:2023bat . As photons from compact stars travel through the scalar field background, their dispersion relation changes due to this interaction, causing the photon wavenumber to change from the point of emission to detection. We study the propagation of pulsar light through the background scalar field.
  • In a binary system of two compact stars, an ultralight scalar particle can mediate a long-range force in addition to the gravitational force between the stars. Various fifth-force experiments can place constraints on such long-range interactions Hook:2017psm ; KumarPoddar:2019jxe .
  • The scalar-induced magnetic field can alter the surface magnetic field of the compact star, which plays a crucial role in determining the energy loss through magnetic dipole radiation Domcke:2023bat .
  • If the source is time-dependent, the scalar field itself can also act as a form of radiation, carrying away energy from the compact star. This leads to a decrease in their spin rate, a process known as spin-down Contopoulos:2005rs ; Lyne:2010ad ; Tauris:2012ex .
The measurements of observables corresponding to the above effects would allow us to constrain the scalar-photon coupling.
In this paper, we do not explore the detailed mechanisms of scalar mass generation, since our focus is not on model building or ultraviolet completion. However, it is known that such ultralight scalars can naturally acquire mass in several theoretical frameworks. These include clockwork mechanism Giudice:2016yja ; Wood:2023lis , Planck-suppressed operators from string theory or quantum gravity Hui:2016ltb ; Hubisz:2024hyz ; Banerjee:2022wzk , and non-perturbative effects such as instantons Kitano:2021fdl . Ultralight scalars also commonly appear in extra-dimensional compactifications Anchordoqui:2023tln , or scale–invariant symmetry breaking scenarios Ferreira:2020fam .
The paper is organized as follows. In Section II, we obtain the scalar field profile due to the scalar-photon interaction outside the compact star. The scalar-induced electric and magnetic fields are calculated in Section III. In Section IV, we derive the modified photon dispersion relation and calculate the modification of the redshift and photon wavenumber in space due to the scalar-photon interaction. The rate of energy loss due to scalar radiation is derived in Section V. In Section VI, we obtain constraints on the strength of scalar-photon interactions based on the searches for a new long-range force in a double pulsar binary, the EM radiation generated by a scalar-induced magnetic field, and pulsar spin-down measurements. Finally, in Section VII we conclude and discuss our results.
We use the system of units with the speed of light in vacuum $c=1$, the reduced Planck constant $\hbar=1$, and the Newton’s gravitational constant $G=1$ throughout the paper, unless stated otherwise.
II Long-range scalar field outside a compact star
A rotating compact star like a NS or a magnetar is a large dipole magnet. In the aligned rotator model (where the magnetic dipole moment is along the rotation axis of the star), the external dipolar magnetic field is given by Goldreich:1969sb ; Shapiro:1983du ; Mohanty:1993nh
[…]
Here, $R$ denotes the radius of the star, $B_{0}$ denotes the magnetic field strength at its surface $(r=R)$, and $\theta$ denotes the polar angle which is measured with respect to the rotation axis of the star. Using the boundary condition that the tangential component of the electric field is continuous at $r=R$ while the normal component of the electric field may be discontinuous across the boundary, the expression for the electric field profile outside the star is Goldreich:1969sb ; Shapiro:1983du ; Mohanty:1993nh
$\mathbf{E}^{\mathrm{out}}_{(r>R)}=-\frac{B_{0}\Omega R^{5}}{r^{4}}\Big[\Big(1-\frac{3}{2}\sin^{2}\theta\Big)\hat{r}+\sin\theta\cos\theta\,\hat{\theta}\Big$]\,, (2)
where $\Omega$ denotes the angular velocity of the star. Using Eqs. 1 and 2, we can estimate the quantity $\frac{1}{2}F_{\mu\nu}F^{\mu\nu}=\mathbf{B}^{2}-\mathbf{E}^{2}$ outside the star as
$\mathbf{B}^{2}-\mathbf{E}^{2}=\frac{B^{2}_{0}R^{6}}{4r^{6}}(3\cos^{2}\theta+1)-\frac{B^{2}_{0}\Omega^{2}R^{10}}{4r^{8}}(5\cos^{4}\theta-2\cos^{2}\theta+1)\,$, (3)
where $F^{\mu\nu}$ denotes the EM stress tensor.
To study the scalar field profile sourced by the EM fields outside a compact star, we write the Lagrangian for a CP-even scalar field interacting with the EM field as
$\mathcal{L}=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}g_{\phi\gamma\gamma}\phi F_{\mu\nu}F^{\mu\nu},$ (4)
where $\phi$ denotes the scalar field and $g_{\phi\gamma\gamma}$ denotes the effective coupling of this scalar field with the EM fields of the star. Note that $g_{\phi\gamma\gamma}$ can have either sign, as determined by the nature of the charged particles running in the triangle loop diagrams responsible for the effective scalar-photon interaction. When the loop contribution is dominated by fermions such as leptons or heavy quarks, the coupling $g_{\phi\gamma\gamma}$ is positive. However, if the loop is dominated by a $W$ boson, the coupling would acquire a negative sign Flambaum:2024zyt .
The equation of motion of the scalar field can be obtained as
$\Box\phi=-g_{\phi\gamma\gamma}(\mathbf{B}^{2}-\mathbf{E}^{2})\,$, (5)
where the d’Alembertian operator is $\Box=\frac{\partial^{2}}{\partial t^{2}}-\nabla^{2}$ in the Minkowski spacetime. Therefore, to have a non-trivial scalar field profile outside the compact star, we must have a nonzero “source charge density” $\rho_{\phi}=g_{\phi\gamma\gamma}(\mathbf{B}^{2}-\mathbf{E}^{2})$ outside the star. Now, for a rotating NS where the angular velocity is not very large, $|\mathbf{B}|\gg|\mathbf{E}|$, and we can neglect the $\mathbf{E}^{2}$ term to write Eq. 5 as
$\Box\phi\approx-g_{\phi\gamma\gamma}\frac{B^{2}_{0}R^{6}}{4r^{6}}(3\cos^{2}\theta+1)\,.$ (6)
Solving Eq. 6 by the Green’s function method in the Schwarzschild background, we obtain the scalar field profile as
$\phi(r)\approx\frac{Q_{\phi}^{\mathrm{eff}}}{r}+\mathcal{O}\Big(\frac{1}{r^{2}}\Big)\,$, (7)
where the effective scalar charge $Q_{\phi}^{\mathrm{eff}}$ is
$Q_{\phi}^{\mathrm{eff}}=\frac{g_{\phi\gamma\gamma}B^{2}_{0}R^{6}}{48M^{3}}\;.$ (8)
Here, $M$ is the mass of the compact star. Note that the scalar-induced effective charge, $Q^{\mathrm{eff}}_{\phi}$, can be either positive or negative, depending on the sign of the coupling $g_{\phi\gamma\gamma}$. The dependence of $Q^{\text{eff}}_{\phi}$ on $M$ in Eq. 8 is a result of general relativistic corrections to the scalar field profile, arising from the Schwarzschild metric. In flat spacetime, the effective charge simplifies to $Q^{\text{eff}}_{\phi}\sim g_{\phi\gamma\gamma}B_{0}^{2}R^{3}$. This form arises from the term $F_{\mu\nu}F^{\mu\nu}\propto|\mathbf{B}|^{2}$ in the scalar source charge density $\rho_{\phi}$, under the assumptions of slow rotation and a magnetically dominated regime $|\mathbf{B}|\gg|\mathbf{E}|$.
The $1/M^{3}$ dependence of $Q_{\phi}^{\mathrm{eff}}$ in Eq. 8 can be motivated as follows. The scalar field profile is obtained by solving Eq. 6, where the $M$- dependence is implicit through the d’Alembertian operator. This operator gives general-relativistic $(M/R)$-corrections to the flat spacetime solution. Given that the flat spacetime solution is $Q^{\mathrm{eff}}_{\phi}\propto g_{\phi\gamma\gamma}B_{0}^{2}R^{3}$ while Eq. 6 forces us to have $Q^{\mathrm{eff}}_{\phi}\propto g_{\phi\gamma\gamma}B_{0}^{2}R^{6}$, dimensional analysis tells us that the $(M/R)$-dependence should be of the form $Q^{\mathrm{eff}}_{\phi}\propto g_{\phi\gamma\gamma}B_{0}^{2}R^{3}(R/M)^{3}$. Explicit calculation also gives the same dependence on $M$.
Note that while deriving the scalar field profile above, the metric inside the star was taken to be Schwarzschild for analytical simplicity. This approximation is justified since replacing the Schwarzschild interior with a full TOV (Tolman-Oppenheimer-Volkoff) solution would modify the scalar charge by only a few percent for a typical NS. One gets $\Delta Q_{\phi}/Q_{\phi}\sim\mathcal{O}(C\,p/\rho)$, where $C\equiv M/R$ denotes the stellar compactness, and $p$ and $\rho$ represent the pressure and density inside the star, respectively. Such a small correction ($\Delta Q_{\phi}/Q_{\phi}\lesssim 10\%$, for typical NS parameters) is well within the astrophysical uncertainties in the stellar radius and magnetic field, making the approximation adequate for the level of accuracy considered in this work.
Equations 7 and 8 indicate that the rotating star has a long-range scalar “hair” associated with a charge $Q_{\phi}^{\mathrm{eff}}$. Though we have obtained these results for a massless scalar, our results would be valid as long as the Compton wavelength of the scalar is greater than the radius of the star, i.e., for $1/m_{\phi}\gtrsim R$, or $m_{\phi}\lesssim 1/R$.
We use GRB 080905A as a benchmark and write Eq. 8 as
$Q_{\phi}^{\mathrm{eff}}=2.4\times 10^{38}\Big(\frac{g_{\phi\gamma\gamma}}{10^{-15}~\mathrm{GeV^{-1}}}\Big)\Big(\frac{B_{0}}{3.93\times 10^{16}~\mathrm{G}}\Big)^{2}\Big(\frac{R}{10~\mathrm{km}}\Big)^{6}\Big(\frac{1.4~M_{\odot}}{M}\Big)^{3}.$ (9)
Note that, for stars with a large angular velocity ($\Omega R\sim{\cal O}(1)$), the electric field outside the star cannot be neglected, and we need to solve the scalar field profile sourced by $\mathbf{B}^{2}-\mathbf{E}^{2}$ instead of only $\mathbf{B}^{2}$. A detailed analysis of this scenario is presented in Appendix A.
III Scalar-induced EM fields from Maxwell’s equations
The interaction of a CP-even scalar with the EM fields of the star modifies Maxwell’s equations for the EM fields in vacuum. We derive the electric and magnetic field equations in a perturbative way by expanding the stress tensor in powers of $g_{\phi\gamma\gamma}$, such that
$F^{\mu\nu}=F^{\mu\nu}_{(0)}+F^{\mu\nu}_{\phi}+\mathcal{O}(g^{2}_{\phi\gamma\gamma})\,$, (10)
where the “(0)(0)” corresponds to any quantity in the limit $g_{\phi\gamma\gamma}=0$. We keep the terms which are linear in $g_{\phi\gamma\gamma}$, and obtain $\partial_{\mu}F^{\mu\nu}_{\phi}=-g_{\phi\gamma\gamma}(\partial_{\mu}\phi)F^{\mu\nu}_{(0)}$ in the absence of source charge and current density of the plasma. This relation gives the expressions for the scalar induced electric $(\mathbf{E}_{\phi})$ and magnetic $(\mathbf{B}_{\phi})$ fields in terms of the background electric $(\mathbf{E}_{(0)})$ and magnetic $(\mathbf{B}_{(0)})$ fields as
$\nabla\cdot\mathbf{E}_{\phi}$ $=$ $-g_{\phi\gamma\gamma}\,\mathbf{E}_{(0)}\cdot\nabla\phi\,$,
$\nabla\times\mathbf{B}_{\phi}$ $=$ $\frac{\partial\mathbf{E}_{\phi}}{\partial t}-g_{\phi\gamma\gamma}\,\nabla\phi\times\mathbf{B}_{(0)}+g_{\phi\gamma\gamma}\left(\frac{\partial\phi}{\partial t}\right)\mathbf{E}_{(0)}\,$, (11)
while the Bianchi identity $\partial_{\mu}\tilde{F}^{\mu\nu}_{\phi}=0$ gives
$\nabla\cdot\mathbf{B}_{\phi}$ $=$ $0\,$,
$\nabla\times\mathbf{E}_{\phi}$ $=$ $-\frac{\partial\mathbf{B}_{\phi}}{\partial t}\,.$ (12)
Note that in the aligned rotator model, the scalar field and the background EM fields do not have any temporal dependence. Further, since the source terms (arising from the background fields) are time-independent, the terms $\partial\mathbf{E}_{\phi}/\partial t$, $\partial\mathbf{B}_{\phi}/\partial t$ and $\partial\phi/\partial t$ also vanish. The scalar-induced magnetic and electric fields are produced due to the interaction of background magnetic (Eq. 1) and electric (Eq. 2) fields with the scalar. For static background EM fields, Eqs. 11 and 12 represent how these background fields are altered due to the interaction with the scalar field.
Combining Eqs. 11 and 12, we obtain the wave equations for the scalar-induced magnetic and electric fields (in the static case) as
[…]
where we neglect terms which appear as two spatial derivatives of $\phi$ (since $\phi$ falls as $1/r$ and the derivatives of $\phi$ will fall even faster). In the limit $\Omega R\ll 1$, we solve Eq. 13 in the Schwarzschild background to obtain the scalar-induced magnetic field as [^4] The form of the scalar-induced magnetic field for large angular velocity is given in Appendix A.
[…]
where we use Eq. 1 for the background magnetic field.
Note that the scalar-induced magnetic field $\mathbf{B}_{\phi}$ in Eq. 14 is actually proportional to $g^{2}_{\phi\gamma\gamma}$, as $Q^{\mathrm{eff}}_{\phi}$ itself is proportional to $g_{\phi\gamma\gamma}$. It falls as $1/r^{2}$ in the radial direction and $1/r$ in the angular direction, as compared to the background magnetic field $\mathbf{B}_{(0)}$ which falls as $1/r^{3}$ both in radial and angular directions. The scalar-induced electric field $\mathbf{E}_{\phi}$ can also be calculated in a similar manner. Since these scalar-induced EM fields scale quadratically with $g_{\phi\gamma\gamma}$, the deviations of the EM fields from their background values, as a result of their interactions with the scalar, are small.
IV EM wave propagation in the background of a long-range scalar field
Maxwell’s equations of electrodynamics for the propagation of light are modified due to the interactions of the CP-even scalar $\phi$ with the EM fields. Consider a situation where light, i.e., an EM wave, is emitted by the source in the presence of the background static EM and scalar fields. In the absence of any other source plasma charge and current densities, the Maxwell’s equations become Domcke:2023bat
$\nabla\cdot\mathbf{E}$ $=$ $-g_{\phi\gamma\gamma}\mathbf{E}\cdot\nabla\phi\,$,
$\nabla\times\mathbf{B}$ $=$ $\frac{\partial\mathbf{E}}{\partial t}-g_{\phi\gamma\gamma}\nabla\phi\times\mathbf{B}\,$,
$\nabla\cdot\mathbf{B}$ $=$ $0\,$,
$\nabla\times\mathbf{E}$ $=$ $-\frac{\partial\mathbf{B}}{\partial t}\,.$ (15)
where the $\mathbf{E}$ and $\mathbf{B}$ are the electric and magnetic fields of the propagating EM wave. In addition, we neglect terms which appear as two spatial derivatives of $\phi$, as earlier. Using Eq. 15, we obtain the equation for the EM wave as
$\begin{split}\Box\mathbf{B}=g_{\phi\gamma\gamma}(\nabla\phi\cdot\nabla)\mathbf{B}\,,\\ \Box\mathbf{E}=g_{\phi\gamma\gamma}(\nabla\phi\cdot\nabla)\mathbf{E}\,.\end{split}$ (16)
We choose the Eikonal ansatz $\mathbf{B}(x,t)=\mathcal{B}\,e^{iS(x,t)}$ for the propagation of light. The phase $S$ defines the frequency and wavenumber of photon along the ray orbit, since $\omega=-\partial S/\partial t$ and $\mathbf{k}=\nabla S$. Eq. 16 implies that in the asymptotically flat spacetime, the dispersion relation of photons is modified due to the scalar field contribution as [^5] Note that, unlike the CP-odd pseudoscalar coupling, the CP-even structure of the source in our case ensures that scalar-photon interactions do not produce any birefringence effects, i.e. the propagation is independent of the photon polarization.
$\omega^{2}=k^{2}-ig_{\phi\gamma\gamma}(\nabla\phi\cdot\mathbf{k})\,$, (17)
where the imaginary component implies that the wavenumber of photon must have a complex value in the presence of a scalar background.
We obtain the expression for the group velocity from Eq. 17 as
$v_{g}=\frac{d\omega}{dk}=\frac{2k-im_{\gamma}}{2\omega(k)},$ (18)
where we define the scalar-induced photon mass $m_{\gamma}=|g_{\phi\gamma\gamma}\nabla\phi|$, assuming the photon propagation is radial. We solve Eq. 17 for $k$ as
$k=k_{R}+ik_{I}=\frac{\sqrt{4\omega^{2}-m_{\gamma}^{2}}}{2}+\frac{im_{\gamma}}{2},$ (19)
where $k_{R}=\sqrt{4\omega^{2}-m_{\gamma}^{2}}/2$ and $k_{I}=m_{\gamma}/2$. Note that $k_{R}$ is a real quantity provided $\omega>m_{\gamma}/2$. Using Eq. 19, we obtain the expression for the group velocity in Eq. 18 as
$v_{g}=\Big(1-\frac{m_{\gamma}^{2}}{4\omega^{2}}\Big)^{\frac{1}{2}}.$ (20)
In the region $\omega\gtrsim m_{\gamma}/2$, the group velocity is subluminal and can be approximated as $v_{g}\approx 1-(m_{\gamma}^{2}/8\omega^{2})$. We observe that the group velocity of the photon remains unchanged at order $\mathcal{O}(g^{2}_{\phi\gamma\gamma})$ since $\nabla\phi\propto g_{\phi\gamma\gamma}$; it receives corrections only at $\mathcal{O}(g^{4}_{\phi\gamma\gamma})$. If there is no scalar-photon coupling, then $m_{\gamma}\rightarrow 0$ and $v_{g}$ becomes unity.
The wavelength of the propagating wave is governed by the real part of the wavenumber, $k_{R}$, in the region $\omega>m_{\gamma}/2$. In this region, $k_{R}\simeq\omega-(m^{2}_{\gamma}/8\omega)$. The contribution of this dispersion relation to the apparent redshift of the photon of wavelength $\lambda$, as measured in the asymptotically flat spacetime (i.e., at the observer point $r_{2}$), is then
$\delta z=\frac{\lambda(r_{2})-\lambda(r_{1})}{\lambda(r_{1})}\approx\frac{k_{R}(r_{1})-k_{R}(r_{2})}{k_{R}(r_{2})}\approx\frac{m^{2}_{\gamma}}{8\omega^{2}}\approx\frac{g^{4}_{\phi\gamma\gamma}B^{4}_{0}R^{8}}{48^{2}\times 8M^{6}\omega^{2}},$ (21)
where, in the last approximations, we consider the relevant magnitude of the apparent redshift and assume that the wave propagates along the direction of $\nabla\phi$, i.e., radially. Here $r_{1}$ represents a location close to the magnetar. We have assumed $m_{\gamma}(r_{2})\approx 0$ as the observer is far away from the magnetar (for example, at the Earth). From Eq. 21, the correction to the photon redshift would be significant if $m_{\gamma}(r_{1})$ is of the same order of magnitude as $\omega$. Moreover, the redshift as measured at different frequencies will be different, an indication of a non-trivial dispersion relation.
The wavelength-dependence of the redshift as indicated in Eq. 21 implies that, if we are able to have measurements of multiple spectral lines from a magnetar (which will yield different redshift values for different photon frequencies) and the redshift of the host galaxy (for an appropriate normalization), we will be able to determine the value of $\delta z(\omega)$ and hence, the value of $g_{\phi\gamma\gamma}$. To present an estimation for the order of magnitude of $\delta z$, we use as the benchmark GRB 080905A Rowlinson:2013ue ; Ricci:2020oyt , which originates from a magnetar. The apparent redshift, or the fractional change in the photon wavenumber, can be expressed using Eq. 21 as
$\delta z=\frac{\Delta k}{k}\sim 10^{-4}\Big(\frac{g_{\phi\gamma\gamma}}{10^{-15}~\mathrm{GeV}^{-1}}\Big)^{4}\Big(\frac{2.1~\mathrm{GHz}}{\omega}\Big)^{2}\Big(\frac{B_{0}}{3.93\times 10^{16}~\mathrm{G}}\Big)^{4}\Big(\frac{R}{10~\mathrm{km}}\Big)^{8}\Big(\frac{1.4~M_{\odot}}{M}\Big)^{6},$ (22)
where we have used Eqs. 7 and 8 to describe the scalar field, and taken $k\approx\omega$ at the leading order. If the redshift measurements have a precision of $\sim 10^{-4}$ Rowlinson:2010jb , we would get a sensitivity of $g_{\phi\gamma\gamma}\sim 10^{-15}$ GeV-1 to the scalar-photon coupling. The shift in the wavenumber would be more pronounced for compact stars that have strong surface magnetic fields, larger dimensions, and emit signals detectable at lower frequencies.
With the advancements in precision atomic clocks, it may be possible to determine the wavelength (or frequency) of a particular spectral line emitted by the magnetar to a precision of $\Delta k/k\sim 10^{-18}$. From Eq. 22, this precision would correspond to a sensitivity of $g_{\phi\gamma\gamma}\sim 10^{-19}~\mathrm{GeV}^{-1}$. This prospective bound has been indicated by a dashed purple line in FIG. 1. The sensitivity can be further enhanced by observing low-frequency photons, which are accessible to radio telescope facilities such as the Square Kilometre Array (SKA) Huege:2016jvc and LOw Frequency ARray (LOFAR) Geyer:2017lbi ; vanHaarlem2013 . The constraint on $g_{\phi\gamma\gamma}$ can be further strengthened through the use of nuclear clocks, entangled clock networks, or high-precision space-based timekeeping systems Kessler:2014sct ; Arakawa:2023gyq ; Fuchs:2024xvc ; PhysRevA.81.030302 ; Yang:2025agr , by employing a large number of atoms, extending the interrogation time, and minimizing environmental noise.
In particular, conventional atomic clocks operate at the standard quantum limit (SQL) with $N$ uncorrelated atoms, the phase evolution scales as $1/\sqrt{N}$. If the atoms are prepared in maximally entangled (fully correlated) states, the ensemble behaves as a single coherent quantum system and the precision can reach the Heisenberg limit (HL), where the phase evolution scales as $1/N$. While these conditions represent near-ideal scenarios, ongoing technological developments are steadily improving the performance of realistic systems.
The corresponding sensitivity curve for a relative wavenumber deviation of $\Delta k/k\sim 10^{-24}$ at a frequency of $\omega\sim 10~\mathrm{MHz}$ (a frequency accessible at the LOFAR telescope) is shown as a green dashed line in FIG. 1
In the regime $\omega>m_{\gamma}/2$, the imaginary part of the wavenumber, $k_{I}$, leads to a damping of the wave amplitude, indicating photon absorption in the scalar field medium. This attenuation manifests in the decay of the photon’s intensity, which is related to the electric field as $I\propto\mathbf{E}^{2}\propto E_{0}^{2}e^{-2k_{I}x}$. Consequently, if a photon with initial intensity $I_{0}$ propagates through an absorptive medium of thickness $x$, the transmitted intensity becomes $I=I_{0}e^{-\alpha x}$, where $\alpha$ represents the absorption coefficient. Within the framework of our analysis, [^6] There could be enhancement in the photon intensity if $(\nabla\phi\cdot\bm{k})$ picks negative sign.
$\alpha=2k_{I}=m_{\gamma}=\frac{1}{x}\ln\Big(\frac{I_{0}}{I}\Big).$ (23)
Therefore, unlike the photon wavelength shift—which appears only at order $\mathcal{O}(g^{4}_{\phi\gamma\gamma})$—the absorption coefficient arises at the order $\mathcal{O}(g^{2}_{\phi\gamma\gamma})$.
The scalar-induced redshift may be distinguished from the possible effects of the gravitational redshift and plasma dispersion. The gravitational redshift is independent of frequency while the scalar-induced redshift goes as $1/\omega^{2}$. Plasma dispersion effects lead to the same $1/\omega^{2}$ dependence; however, they are accompanied by Faraday rotation which is absent in the scalar-induced case.
V Scalar field radiation from an isolated compact star
In the situations considered in previous sections, the scalar field was coupled to a static source, resulting in a long-range $1/r$ scalar field profile outside the magnetized star. Since a static source does not emit scalar radiation, no scalar field radiation originates from the star in these situations. To investigate the impact of scalar radiation on pulsar spin-down, we now consider a scenario where the EM fields of the magnetar are oscillating with time. We consider a skewed rotator model, where the magnetic moment axis of the star makes an angle $\alpha$ with its rotation axis. In this model, the magnetic field at any space-time point can be written as cite-key
[…]
where the magnetic colatitude $\theta_{m}$ is the angle between the magnetic moment axis and the line of sight. It is expressed as $\cos\theta_{m}=\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos(\Omega t-\varphi)$, where $\alpha$ is the angle between the rotational axis and the magnetic moment axis, and $\theta$ is the angle between the rotational axis and the line of sight. In the limit $\alpha\rightarrow 0$, Eq. 24 reduces to Eq. 1. Thus, for radiation, one needs a non-zero $\alpha$.
The source charge density for $\phi$ may be written as $\rho_{\phi}(\mathbf{r},t)=g_{\phi\gamma\gamma}(\mathbf{B}^{2}-\mathbf{E}^{2})$, where Eq. 24 gives
$\mathbf{B}^{2}-\mathbf{E}^{2}\approx-\frac{3}{2}\frac{B_{0}^{2}R^{6}}{r^{6}}\Big[\sin\alpha\cos\alpha\sin\theta\cos\theta\cos(\Omega t-\phi)+\frac{1}{2}\sin^{2}\alpha\sin^{2}\theta\cos^{2}(\Omega t-\phi)\Big],$ (25)
and we omit terms involving $\Omega R\ll 1$, and hence, $\mathbf{E^{2}}$. We also remove the time-independent terms which do not contribute to the radiation.
In Eq. 25, the term $\mathbf{B}^{2}-\mathbf{E}^{2}\propto\sin\theta\cos\theta$ exhibits a quadrupolar angular dependence, while the $\sin^{2}\theta$ component contains both a time-independent DC part and a quadrupolar contribution. Consequently, the first term leads to quadrupolar radiation at the rotation frequency $\Omega$, whereas the second term produces radiation at twice the frequency, $2\Omega$, in addition to the non-radiative DC component. Therefore, unlike the pseudoscalar axion case Khelashvili:2024sup , no scalar dipole radiation arises in this scenario.
The corresponding source charge densities for radiation at $\Omega$ and $2\Omega$ can therefore be written as
$\rho_{+\Omega}(\mathbf{r})=\frac{3}{2}g_{\phi\gamma\gamma}\mu^{2}\sin(2\alpha)\sqrt{\frac{8\pi}{15}}\frac{Y_{2,-1}(\theta,\phi)}{r^{6}},~~~\rho_{+2\Omega}(\mathbf{r})=-\frac{3}{4}g_{\phi\gamma\gamma}\mu^{2}\sin^{2}\alpha\sqrt{\frac{32\pi}{15}}\frac{Y_{2,-2}(\theta,\phi)}{r^{6}},$ (26)
where $\mu=B_{0}R^{3}/2$ and $Y_{l,m}(\theta,\phi)$ are spherical harmonics.
The scalar field satisfies the wave equation
$(\nabla^{2}+k^{2})\phi_{\omega}(\mathbf{r})=-\rho_{\omega}(\mathbf{r}),~~~k=\sqrt{\omega^{2}-m^{2}_{\phi}},$ (27)
whose solution, using the retarded Green’s function, is
$\phi_{\omega}(\mathbf{r})=\int d^{3}r^{\prime}\frac{e^{ik|\mathbf{r}-\mathbf{r^{\prime}}|}}{4\pi|\mathbf{r}-\mathbf{r^{\prime}}|}\rho_{\omega}(\mathbf{r^{\prime}}).$ (28)
Defining the multipole moments as
$Q_{lm}(\omega)=\int d^{3}r^{\prime}\rho_{\omega}(\mathbf{r}^{\prime}){r^{\prime}}^{l}Y_{l,m}^{*}(\hat{r}^{\prime}),$ (29)
the scalar field in the far-field and long-wavelength limit becomes
$\phi_{\omega}(\mathbf{r})\simeq\frac{e^{ikr}}{r}\sum_{lm}i^{l}\frac{k^{l}}{(2l+1)!!}Y_{l,m}(\hat{r})Q_{lm}(\omega).$ (30)
For a real scalar field, the radial energy flux is
$T^{0r}=\dot{\phi}\partial_{r}\phi,~~~\langle S_{r}\rangle=\langle T^{0r}\rangle=\frac{\omega k}{2}|\phi_{\omega}(\mathbf{r})|^{2},$ (31)
where the angle brackets denote time-averaging over one period. The total power radiated in scalar waves is then
$P_{\omega}=\int d\Omega r^{2}\langle S_{r}\rangle=\frac{\omega k}{2}\sum_{lm}\Big(\frac{k^{l}}{(2l+1)!!}\Big)^{2}|Q_{lm}(\omega)|^{2}.$ (32)
Hence, the scalar quadrupole radiation power at frequencies $\Omega$ and $2\Omega$ are given by
$P_{\Omega}\simeq\frac{1}{80}g^{2}_{\phi\gamma\gamma}B^{4}_{0}R^{10}\Omega^{6}\Big(1-\frac{m^{2}_{\phi}}{\Omega^{2}}\Big)^{5/2}\sin^{2}2\alpha,~~~P_{2\Omega}\simeq\frac{1}{100}g^{2}_{\phi\gamma\gamma}B^{4}_{0}R^{10}\Omega^{6}\Big(1-\frac{m^{2}_{\phi}}{4\Omega^{2}}\Big)^{5/2}\sin^{4}\alpha.$ (33)
Therefore, scalar radiation at frequency $\Omega$ is allowed when $m_{\phi}\lesssim\Omega$, while radiation at $2\Omega$ occurs only if $m_{\phi}\lesssim 2\Omega$. Also, for a typical pulsar, the misalignment angle is small and $P_{\Omega}$ dominates over $P_{2\Omega}$. Scalar radiation can occur only when the pulsar’s magnetic axis is tilted with respect to its rotation axis. In the aligned-rotator limit $\alpha\to 0$, where the two axes coincide, the scalar radiation vanishes.
VI Constraints from Observations
In this section, we employ the results obtained in the previous sections and attain constraints on the scalar-photon coupling based on various observations related to pulsars and magnetars. The constraints may originate from the bounds on the magnitude of a new long-range force in double pulsar binaries, as well as from the measurements of the radiation from a compact star and of its spin-down luminosity.
The most stringent current bounds on the scalar-photon coupling arise from the studies searching for variation in the fine-structure constant caused by the interaction between photons and the scalar field. These bounds are obtained assuming that the scalar field is responsible for the entire DM in the universe. The Holometer bound Aiello:2021wlp is obtained by studying the variation of the fine structure constant $\alpha$ with the cross-correlated data of the Fermilab Holometer instrument. The Cs/Cav result is obtained from the study of the variation of $\alpha$ with the optical spectroscopy apparatus Tretiak:2022ndx . The GEO 600 bound Vermeulen:2021epa is obtained by doing spectral analysis of the strain data of the GEO 600 interferometer. The LIGO bound Fukusumi:2023kqd ; Gottel:2024cfj is obtained from the LIGO-Virgo data, based on studying the variation of $\alpha$. The thin vertical gray-shaded region is excluded by AURIGA Branca:2016rez , where the bound is obtained by studying the oscillation of cryogenic resonant mass AURIGA detector due to the scalar DM. The H/Quartz/Sapphire Campbell:2020fvq bound is obtained from the search for frequency modulation due to oscillating DM interaction in the frequency-stable oscillators such as hydrogen maser atomic oscillator, bulk acoustic wave quartz oscillator, and cryogenic sapphire oscillator. The Dy/Quartz Zhang:2022ewz bound is obtained by comparing the frequency of a quartz oscillator to the hyperfine and electronic transitions of $^{87}\mathrm{Rb}$ and $^{164}\mathrm{Dy}$, respectively, due to effect of time-oscillating DM. The bound for Dynamical Decoupling (DD) Aharony:2019iad is obtained from the non-observation of variation of $\alpha$ due to the oscillating scalar DM in an atomic optical transition. The $I_{2}$ bound Oswald:2021vtc is obtained by studying the oscillations of $\alpha$ induced by the ultralight scalar DM and their effect on the Iodine molecular spectroscopy. The DAMNED bound Savalle:2020vgz is obtained from the search for DM with an optical cavity and unequal delay interferometer. The parameter space for $g_{\phi\gamma\gamma}$ is also constrained by the optical and atomic clock studies for the search of DM, such as PTB Filzinger:2023zrs , Sr/Si Kennedy:2020bac , Rb/Cs Hees:2016gop , Dy/Dy VanTilburg:2015oza , BACON BACON:2020ubh , $\mathrm{Yb^{+}/Sr}$ Sherrill:2023zah . These bounds have been shown with various shades of gray in FIG. 1. However, since the scalar in our scenario need not play the role of DM, these bounds are not directly applicable for our scenario.
On the other hand, there are bounds that do not need the scalar to be the DM. The astrophysical bounds from globular clusters Dolan:2022kul are obtained by calculating the ratio of the energy losses due to the scalar in the asymptotic giant branch to the horizontal branch stars. The bounds from the Eöt-Wash experiment Hees:2018fpg , fifth force experiments Fischbach:1996eq ; Adelberger:2003zx ; Konopliv2011 ; Fienga:2023ocw and MICROSCOPE experiment Berge:2017ovy are obtained from the precision tests of Einstein’s equivalence principle, using laboratory measurements or observations in solar system. These bounds are relevant for comparison and complementarity with our work.
VI.1 Search for a new long-range force in a double pulsar binary
Figure 1: Bounds on $g_{\phi\gamma\gamma}$ derived from the measurements of the electromagnetic radiation by a scalar-induced magnetic field (blue shaded region), and pulsar spin-down caused by scalar radiation (red shaded region). The prospective bound from possible future constraints on a new long-range force from a pulsar binary pair with large surface magnetic fields $B_{0}$ is shown as a magenta dashed line, while the prospective bound from the measurement of the photon wavenumber using a future precision atomic clock ($\Delta k/k\sim 10^{-18},\omega=2.1$ GHz) is shown as a purple dashed line. The sensitivity obtainable with $\Delta k/k\sim 10^{-24}$ at $\omega\sim 10$ MHz has also been shown with a green dashed line. The existing constraints are shown as gray-shaded areas.
We have seen in Section II that the scalar field interaction with the EM fields of a compact star induces a scalar charge on the compact star. For a system of two compact stars in a binary, this would lead to a scalar-mediated long-range force that has the same spatial dependence, $1/r^{2}$, as the gravitational force between the two stars. The ratio of the long-range force to the gravitational force is
$\eta=\frac{Q^{\mathrm{eff}}_{1}Q^{\mathrm{eff}}_{2}}{4\pi GM_{1}M_{2}}\approx\frac{g^{2}_{\phi\gamma\gamma}B^{2}_{01}B^{2}_{02}R^{6}_{1}R^{6}_{2}}{(48)^{2}\times{4\pi}G^{7}M_{1}^{4}M^{4}_{2}}\;,$ (34)
where we use the expression for the scalar charge as given in Eq. 8 and write the Newton’s constant $G$ explicitly. Here $B_{01}$, $B_{02}$ are the surface magnetic fields of the two stars in a binary, and $M_{1}$, $M_{2}$ are their masses, respectively, assuming the two stars to have equal radii $(R_{1}=R_{2})$.
As a concrete example to demonstrate how the scalar-mediated force may be constrained, we consider the double pulsar binary system PSR J0737-3039 Kramer:2006nb ; Kramer2008 . The surface magnetic fields of the two pulsars are $B_{01}\sim 6.3\times 10^{9}~\mathrm{G}$ and $B_{02}\sim 1.2\times 10^{12}~\mathrm{G}$ Lorimer:2007xw . Their masses are $M_{1}=1.3381\pm 0.0007~M_{\odot}$ and $M_{2}=1.2489\pm 0.0007~M_{\odot}$ Kramer:2006nb . We take $R_{1}=R_{2}=10$ km for an estimation. This gives $\eta\sim(1.6\times 10^{7})\ g^{2}_{\phi\gamma\gamma}~\mathrm{GeV}^{2}$. The measurement/bound on $\eta$ can then be translated to the measurement/bound on $g_{\phi\gamma\gamma}$.
However, since the gravitational and the scalar-induced long-range force have the same spatial dependence to the leading order, it would be difficult to separate their contributions by simply measuring the attractive force between them. Indeed, a change in the magnitude of the force could be approximately mimicked by a change in the measured values of the masses. This quasi-no-go situation may be circumvented if the values of the masses are determined independently by some means other than the gravitational measurements, or we have access to a third body that is gravitationally bound with the binary but has a magnetic field much different than the two compact stars.
The distance between the two stars in PSR J0737-3039 is $a=8.8\times 10^{5}~\mathrm{km}$, which is known to $\sim 0.05\%$. If indeed the masses of the stars in PSR J0737-3039 were also known to a precision of $\sim 0.05\%$ as the current measurements suggest, the precision in the prediction of the gravitational force would be $\sim 0.1\%$ and hence the observations would be sensitive to $\eta\sim 10^{-3}$. Not finding a deviation at this level would put a bound of $g_{\phi\gamma\gamma}\lesssim 8\times 10^{-6}~\mathrm{GeV}^{-1}$. Note that this bound would be applicable only when the range of the scalar-induced force is more than the distance between the two compact stars, i.e. $1/m_{\phi}\gtrsim a$ or $m_{\phi}\lesssim 1/a$.
While our ability to obtain a concrete bound on $g_{\phi\gamma\gamma}$ at this stage is limited by the lack of available information about the masses of the stars through non-gravitational means or from a third gravitationally-bound body, future observations may locate a system where these conditions are fulfilled. Since the scalar charge is proportional to the square of the magnetic field strength, larger values of $B_{0}$ will give better constraints on $g_{\phi\gamma\gamma}$. The constraint can be significantly improved for binary magnetar systems because of the larger magnetic fields. So far, no magnetar binary system has been detected. However, future experiments and observations with better sensitivity can explore this possibility Popov:2020ken ; Chrimes:2022imi ; Sherman:2024asw ; Chandrasekhar:1953zz ; Sinha:2010fm .
For a benchmark, consider a binary system consisting of two magnetars, each with a surface magnetic field of $B_{0}\sim 10^{16}~\mathrm{G}$, separated by the same distance as the components of PSR J0737-3039, and having identical masses and radii to the stars in that system. Let us also assume that the masses of the compact stars are measured with a precision of $0.05\%$, so that the measurements are sensitive to $\eta\sim 10^{-3}$. Under these conditions, the projected constraint on the scalar-photon coupling can be obtained as $g_{\phi\gamma\gamma}\lesssim 6\times 10^{-16}~\mathrm{GeV}^{-1}$. If the magnetic fields were $10^{17}$ G each, the corresponding bound will be $g_{\phi\gamma\gamma}\lesssim 6\times 10^{-18}~\mathrm{GeV}^{-1}$. These prospective bounds have been given in TABLE 1. The last bound is represented by the magenta dashed line in FIG. 1.
Search for a new long-range force
Limits PSR J0737-3039 $B_{01,02}\sim 10^{16}~\mathrm{G}$ $B_{01,02}\sim 10^{17}~\mathrm{G}$
$g_{\phi\gamma\gamma}$ $\lesssim 8\times 10^{-6}~\mathrm{GeV}^{-1}$ $\lesssim 6\times 10^{-16}~\mathrm{GeV}^{-1}$ $\lesssim 6\times 10^{-18}~\mathrm{GeV}^{-1}$
Table 1: Summary of the prospective bounds on the scalar-photon coupling. For compact stars separated by $a=8.8\times 10^{5}~\mathrm{km}$, these limits are valid for $m_{\phi}\lesssim 1/a=2.2\times 10^{-16}$ eV.
Precise measurements of post-Keplerian parameters such as the mass ratio, orbital period decay, orbital inclination, periastron advance, Einstein delay, and Shapiro delay enable the determination of individual pulsar masses. Currently, pulsar mass measurements can reach accuracies of about $\mathcal{O}(10^{-3}-10^{-4})$. However, upcoming radio telescopes like MeerKAT Hu:2020ubl and SKA Huege:2016jvc , along with future space-based GW observatories like LISA Thrane:2019lwv , are expected to push this precision further to an unprecedented level of $\mathcal{O}(10^{-6}-10^{-7})$. Specifically, projected measurements of the masses of Pulsars A and B in the double pulsar system PSR J0737–3039A/B by MeerKAT are expected to reach precisions of $\mathcal{O}(10^{-5})$ and $\mathcal{O}(10^{-6})$, respectively Hu:2022jiq . Moreover, MeerKAT is anticipated to measure the moment of inertia of Pulsar A with a precision of about $11\%$. This measurement will provide important constraints on the NS EOS, thereby refining the mass estimates of the pulsars even further Hu:2020ubl .
VI.2 Electromagnetic radiation due to the scalar-induced magnetic field
As discussed in section III, the scalar field interaction with the EM fields of a compact star gives rise to a scalar-induced long-range magnetic field $\mathbf{B}_{\phi}$. If these fields are time-dependent, they result in radiated power, which decreases the rotational energy $E=I\Omega^{2}/2$ of the star, where $I$ is its moment of inertia. The loss of rotational energy results in a decrease in $\Omega$ and hence an increase in the time period of rotation $P$ of the star:
$\frac{dE}{dt}=\frac{d}{dt}\left(\frac{I}{2}\Omega^{2}\right)\approx I\,\Omega\,\dot{\Omega}=I\frac{(2\pi)^{2}}{P^{3}}\dot{P}\;.$ (35)
We conservatively assume that the energy loss is entirely due to the magnetic dipole radiation. The rate of energy released by the magnetic dipole radiation is
$\frac{dE}{dt}\Big|_{\rm magnetic~dipole}=\frac{2}{3}(B_{0}R^{3}\sin\alpha)^{2}\Omega^{4}=\frac{2(2\pi)^{4}}{3}\left(\frac{B_{0}R^{3}\sin\alpha}{P^{2}}\right)^{2}\,.$ (36)
From Eqs. 35 and 36, we get
$B_{0}\sin\alpha=\Big(\frac{3I}{8\pi^{2}R^{6}}\Big)^{\frac{1}{2}}(P\dot{P})^{\frac{1}{2}}\;.$ (37)
The surface magnetic field of the compact star $B_{0}$ has contributions from the standard EM fields and the scalar-induced magnetic field $\mathbf{B}_{\phi}|_{r=R}$ as given in Eq. 14. Taking this into account, the corresponding bound on $g_{\phi\gamma\gamma}$ can be obtained.
We use three sources for our analysis: the Crab pulsar Lyne1993 ; Bejger:2002ty ; Manchester:2004bp ; Philippov:2013aha , the Soft Gamma Repeater SGR 1806-20 sp2010 ; Younes:2015hsa , and GRB 080905A Rowlinson:2010jb ; Rowlinson:2013ue . In TABLE 2, we summarize the input parameters of these compact stars such as their spin period $P$, the period derivative $\dot{P}$, the surface magnetic field $B_{0}$, the radius of the compact star $R$, the inclination angle $\alpha$, and the bounds obtained by us on the scalar-photon coupling. Note that these bounds are valid when the range of the scalar field is more than the radius of the star, i.e. $1/m_{\phi}\gtrsim R$ or $m_{\phi}\lesssim 1/R$.
Search for a scalar induced magnetic field
Crab pulsar SGR 1806-20 GRB 080905A
$P$ $33~\mathrm{ms}$ Lyne:2014qqa $7.468~\mathrm{s}$ Marsden:2001tw $9.80~\mathrm{ms}$ Rowlinson:2013ue
$\dot{P}$ $4.20\times 10^{-13}~\mathrm{s~s^{-1}}$ Lyne:2014qqa $115.7\times 10^{-12}~\mathrm{s~s^{-1}}$ Marsden:2001tw $1.86\times 10^{-7}~\mathrm{s~s^{-1}}$Rowlinson:2013ue
$B_{0}$ $(6.9-8.5)\times 10^{12}~\mathrm{G}$ Khelashvili:2024sup $(8-25)\times 10^{14}~\mathrm{G}$ MEREGHETTI20111317 $(27.1-49.5)\times 10^{15}~\mathrm{G}$ Rowlinson:2013ue
$R$ $14~\mathrm{km}$ Khelashvili:2024sup $10~\mathrm{km}$ Marsden:2001tw $10~\mathrm{km}$ Rowlinson:2013ue
$\alpha$ $70^{\circ}$ Kou:2015oma $70^{\circ}$ Salmonson:2006jk $23^{\circ}$ Rowlinson:2010jb
$g_{\phi\gamma\gamma}$ $\lesssim 6\times 10^{-15}~\mathrm{GeV}^{-1}$ $\lesssim 10^{-17}~\mathrm{GeV}^{-1}$ $\lesssim 5\times 10^{-18}~\mathrm{GeV}^{-1}$
$m_{\phi}$ $\lesssim 1.4\times 10^{-11}~\mathrm{eV}$ $\lesssim 2\times 10^{-11}~\mathrm{eV}$ $\lesssim 2\times 10^{-11}~\mathrm{eV}$
Table 2: Input Parameters for the candidate compact stars and bounds obtained on the scalar-photon coupling $g_{\phi\gamma\gamma}$, from the search for a scalar-induced magnetic field. These bounds are valid for the ranges of $m_{\phi}$ as shown.
In order to put bounds on $g_{\phi\gamma\gamma}$, we express the observed surface magnetic field as
$B_{\mathrm{surf}}^{2}\simeq B_{0}^{2}+B_{\phi}^{2}(g_{\phi\gamma\gamma},B_{0},M,R)\;.$ (38)
We further impose the condition that the additional magnetic contribution $B_{\phi}$ does not exceed the $1\sigma$ measurement uncertainty in $B_{0}$. Uncertainties in the stellar mass and radius are treated as nuisance parameters and propagated through $B_{\phi}$.
The strongest bound on the coupling at the polar cap is obtained from GRB 080905A, which is valid for $m_{\phi}\lesssim 2\times 10^{-11}$ eV. This bound is eight orders of magnitude stronger than the astrophysical bound obtained using globular clusters, and is stronger than the current fifth-force bound for $m_{\phi}\lesssim 1.5\times 10^{-22}~\mathrm{eV}$. This constraint is depicted in FIG. 1 by the blue-shaded region. The bounds for Crab pulsar and SGR 1806-20 are weaker than GRB 080905A and we do not show them in the figure.
The reason why our method yields stronger bounds than that from the fifth-force measurements at ultralight masses is as follows. The fifth-force experiments investigate the derivative of a generic Yukawa potential across small and large length scales to identify any deviations from standard gravity. In the very long-range limit, the mass of the mediator approaches zero, making the Yukawa potential indistinguishable from the standard Newtonian potential. In our method, the constraints on the coupling are valid as long as the scalar mass is smaller than the relevant inverse-distance scale in the observed system: the inverse of the distance between the binaries, the inverse radius of the compact star, or the spin frequency of the compact star. Therefore, our results stay valid as $m_{\phi}$ goes to zero.
VI.3 Spin-down of compact stars due to scalar radiation
As discussed in the preceding subsection, the rotational energy of a compact star, and thus its spin, decreases due to the EM radiation. The gravitational wave radiation would also contribute, but its contribution would be negligible. The scalar radiation may also contribute to this spin-down, which can be measured using $\dot{P}$, i.e., the rate of increase of the spin period $P$. The spin-down luminosity of the star is the rate of loss of its rotational energy, $dE/dt$, as given in Eq. 35.
To constrain the scalar-photon coupling, we assume that the total energy loss rate of the pulsar receives contributions from both standard mechanisms and the scalar radiation given by Eq. 33. Since the inclination angle $\alpha$ is generally small, scalar emission at frequencies $m_{\phi}\lesssim\Omega$ is expected to dominate. For a conservative estimate, we set $\sin^{2}2\alpha=1$. It is worth noting that this mass range is narrower than the condition $m_{\phi}\lesssim 1/R$ discussed in the previous subsection, as for Crab pulsar $\Omega\sim 10^{-13}~\mathrm{eV}$ while $1/R\sim 10^{-11}~\mathrm{eV}$.
The spin-down bounds are derived by comparing the observed spin-down luminosity $\dot{E}_{\mathrm{obs}}$ (Eq. 35), with the theoretical sum of the standard electromagnetic loss $\dot{E}_{\mathrm{EM}}$ (Eq. 36) and the scalar-radiation power $\dot{E}_{\phi}$ (Eq. 33). Using the measured $P,\dot{P}$, and $B_{0}$ from the vacuum dipole model, we require $\dot{E}_{\mathrm{EM}}+\dot{E}_{\phi}\leq\dot{E}_{\mathrm{obs}}$. We obtain a bound on $g_{\phi\gamma\gamma}$ by requiring that the scalar-induced energy loss $\dot{E}_{\phi}$ remains within the $1\sigma$ measurement uncertainty of the observed electromagnetic spindown power $\dot{E}_{\mathrm{EM}}$.
In TABLE 3, we summarize the values of spin-down luminosity, the bounds on scalar-photon coupling and scalar mass from the spin-down of Crab pulsar, SGR 1806-20, and GRB 080905A. The strongest bound on the coupling is obtained from the spin-down of GRB 080905A.
Spin-down of compact stars
Crab pulsar SGR 1806-20 GRB 080905A
$dE/dt$ $(4.49-4.51)\times 10^{38}~\mathrm{erg/s}$Khelashvili:2024sup $(0.5-1.4)\times 10^{36}~\mathrm{erg/s}$ MEREGHETTI20111317 $(0.7-3.8)\times 10^{48}~\mathrm{erg/s}$ Rowlinson:2013ue
$g_{\phi\gamma\gamma}$ $\lesssim 2\times 10^{-11}~\mathrm{GeV}^{-1}$ $\lesssim 7\times 10^{-9}~\mathrm{GeV}^{-1}$ $\lesssim 2\times 10^{-14}~\mathrm{GeV}^{-1}$
$m_{\phi}$ $\lesssim 1.2\times 10^{-13}~\mathrm{eV}$ $\lesssim 5.5\times 10^{-16}~\mathrm{eV}$ $\lesssim 4.2\times 10^{-13}~\mathrm{eV}$
Table 3: Input Parameters for the candidate compact stars and bounds obtained on the scalar-photon coupling $g_{\phi\gamma\gamma}$ for a range of scalar masses, from the measurements of spin-down.
In FIG. 1, the constraint obtained from the spin-down of GRB 080905A is shown in the red-shaded region, while the bounds from the Crab pulsar and SGR 1806-20 are not depicted, as they are weaker in comparison. The constraints from Crab pulsar and GRB 080905A are stronger than the astrophysical bounds from globular clusters. However, they are still weaker than the bounds from Eöt-Wash Hees:2018fpg and MICROSCOPE Berge:2017ovy experiments. More sensitive pulsar spin-down measurements could lead to bounds exceeding those set by the laboratory tests for the equivalence principle.
VII Conclusions and Discussions
Ultralight scalar particles can couple with the time-independent electric and magnetic fields of a compact star, which would result in a long-range scalar field around the star with a spatial dependence $\phi\sim 1/r$. Several laboratory and astrophysical measurements, such as the tests of the equivalence principle from Eöt-Wash experiment, the fifth force experiments, and measurements in atomic spectroscopy, yield stringent constraints on the EM couplings of these scalars. In this paper, we propose and analyze multiple ways of constraining these couplings using observations of pulsars, magnetars and double pulsar binaries.
The $(\sim 1/r)$ spatial dependence of the scalar field differs from the $(\sim 1/r^{2})$ spatial dependence of the pseudoscalar axions that may couple to the EM field. Due to this spatial dependence, the scalar scenario may be considered equivalent to having a scalar charge $Q_{\phi}$ on the star, giving rise to a Coulomb-like long-range potential Mohanty:1993nh . The effects of this long-range scalar “hair” would be significant till a distance $r\sim 1/m_{\phi}$ outside the star, and would affect multiple observations.
We work in the context of a theory with an effective scalar-photon coupling $g_{\phi\gamma\gamma}$. Such a coupling would typically arise from a UV theory with a mass scale $\Lambda$. In our work, $g_{\phi\gamma\gamma}$, and hence this $\Lambda$, are effective parameters and we are agnostic about their origin. The bounds obtained on $g_{\phi\gamma\gamma}$ can be translated to the bounds on $\Lambda$ in the context of a UV-complete theory, but it is beyond the scope of our work. The mass of the ultralight scalar, $m_{\phi}$, may also be related to $\Lambda$ in the context of a UV-complete theory. However, in the phenomenological work of ours, it is simply a free mass parameter.
The interaction of the scalar with the EM fields modifies Maxwell’s EM equations and gives rise to scalar-induced electric and magnetic fields. The dispersion relation of the EM radiation (photon) emitted by the star is also modified during its propagation through the long-range scalar field. This would result in a wavelength-dependent apparent redshift of photons emitted by the star. The measurement of this wavelength dependence, combined with the knowledge of the redshift of host galaxy, can lead to the determination of $g_{\phi\gamma\gamma}$. With the currently possible precision on redshifts ($\delta z\sim 10^{-4}$), one can be sensitive to $g_{\phi\gamma\gamma}\sim 10^{-15}$ GeV-1, using the benchmark GRB 080905A. With future precision atomic clocks that may be able to measure a specific spectral line from the magnetar with a precision of $\delta z\sim\Delta k/k\sim 10^{-18}$, the sensitivity to $g_{\phi\gamma\gamma}\sim 10^{-19}~\mathrm{GeV}^{-1}$ may be obtained. With precision measurements of low-frequency photon wavenumbers using entangled quantum clock networks, the sensitivity to the scalar-photon coupling $g_{\phi\gamma\gamma}$ would be competitive with the existing bounds.
In a binary pulsar system where both of the compact stars give rise to scalar fields, the stars experience a long-range scalar-mediated force in addition to gravity, arising from the interaction between the scalar fields and EM fields sourced by the two stars. The scalar-mediated force may be mimicked by a change in the masses of the two stars. However, if independent information about the masses of the two stars and the distance between them is available – for example, from their interaction with a third body gravitationally bound to them but without a large EM field – then it would be possible to detect this force or to constrain its value. Using the parameters of the pulsar binary system PSR J0737-3039, if the masses of the two stars and the distance between them is known to a precision of $0.05\%$, we find that a constraint of $g_{\phi\gamma\gamma}\lesssim 8\times 10^{-6}~\mathrm{GeV}^{-1}$ on the scalar-photon coupling may be obtained. This constraint would be valid for a scalar mass of $m_{\phi}\lesssim 2.2\times 10^{-16}~\mathrm{eV}$ so that the range of the force is more than the distance between the two stars in this binary system. This bound is weaker than that from the fifth force measurements by several orders of magnitude. However, it is inversely proportional to the square of the magnetic field at the surface of each star, and future measurements of a binary magnetar system with a high magnetic field $(B_{0}\sim 10^{17}~\mathrm{G})$ could improve it to $g_{\phi\gamma\gamma}\lesssim 10^{-18}~\mathrm{GeV}^{-1}$ if the masses of the stars and the distance between them are known to $\sim 0.05\%$.
If the background EM fields are time-dependent, the scalar-induced EM fields, through their radiation, would also carry away additional energy from the source star. This would result in a decrease in the rotational energy of the star, and a consequent increase in its spin period. The surface magnetic field of a compact star may be predicted from the measurements of the spin period and its derivative, which can be well measured from radio and X-ray observations. Indeed, a bound of $g_{\phi\gamma\gamma}\lesssim 5\times 10^{-18}~\mathrm{GeV}^{-1}$ may be obtained from the observations of GRB 080905A. This bound is valid for the $m_{\phi}\lesssim 2\times 10^{-11}~\mathrm{eV}$, which ensures that the range of the scalar field is more than the size of the star.
If the background EM fields are time-dependent, the scalar field will also be time-dependent and hence will radiate. This scalar radiation will lead to an additional spin-down of the star. The spin-down luminosity, or the rate of change of rotational energy of the star, is a measurable quantity and can be obtained from the measurements of the spin period and its derivative. The scalar spin-down luminosity increases with increasing surface magnetic field, radius, and spin frequency of the star. We analyze the data on the Crab pulsar, SGR 1806-20, and GRB 080905A, and obtain the strongest bound on the scalar-photon coupling from the measurement of the spin-down luminosity of GRB 080905A as $g_{\phi\gamma\gamma}\lesssim 2\times 10^{-14}~\mathrm{GeV}^{-1}$ for $m_{\phi}\lesssim 4.2\times 10^{-13}~\mathrm{eV}$.
The constraints discussed here from various observations can be further improved with enhanced sensitivity of detection and by focusing on stars with high surface magnetic fields, larger radii, and higher spin frequencies. The strongest constraint on the scalar-photon coupling comes from the measurements of the rate of change of the spin period due to the EM radiation. Note that the constraints presented on $g_{\phi\gamma\gamma}$ are actually on its magnitude. The expressions for the scalar-induced apparent redshift of photons, the scalar-induced magnetic field, the energy loss rate from pulsars due to scalar radiation, and the scalar-mediated fifth-force between two pulsars all depend on even powers of the coupling $g_{\phi\gamma\gamma}$. As a result, the constraints derived on $g_{\phi\gamma\gamma}$ are insensitive to the sign of the coupling.
Our analysis of electromagnetic radiation due to scalar-induced magnetic field from GRB080905A yields the most stringent astrophysical constraint on $g_{\phi\gamma\gamma}$ to date, improving the previous astrophysical bound from globular clusters by eight orders of magnitude. At extremely low masses, i.e. for $m_{\phi}\lesssim 10^{-22}~\mathrm{eV}$, our bound are stronger than that from the fifth-force experiments. The existing bounds from the atomic clock experiments assume the scalar to be the DM, and hence do not directly apply to our scenario. The MICROSCOPE bound, obtained through the measurement of the variation of nuclear binding energy induced by the change in the fine structure constant, is stronger. However, our bounds are obtained from astrophysical observations, without involving detailed nuclear physics considerations, and are therefore complementary to this bound. In future, developments such as highly sensitive nuclear or space-based entangled clock systems distributed over large baselines, combined with the detection of low frequency photon signals by instruments like LOFAR and SKA, could enable precise measurements of deviations in the photon wavenumber. This would potentially yield limits that are competitive with those from current equivalence principle experiments.
Ultralight scalar particles arise in many theoretical models and motivate diverse experimental searches, regardless of their contribution to DM. We have explored their observable implications within a minimal framework, characterized solely by the scalar mass $m_{\phi}$ and its coupling to photons $g_{\phi\gamma\gamma}$, without invoking hidden-sector assumptions. We point out for the first time that such scalars generate a long-range monopole field outside magnetized stars, and show that, unlike axions, they do not induce birefringence. Future precision measurements of pulsars and magnetars could pave the way for a better understanding of the physics of ultralight scalars, in addition to the astrophysics of these compact objects.
Acknowledgements
The authors would like to thank Shadab Alam, Sumanta Chakraborty, Girish Kulkarni, Jamie Mcdonald, Arunava Mukherjee and Nicholas Rodd for useful discussions. The authors would also like to thank the anonymous referee for their fruitful comments and suggestions. T. K. P. would also like to thank the Galileo Galilei Institute for Theoretical Physics and ICTP for the hospitality and the INFN for partial support during the completion and finalization of this work, and COST Actions COSMIC WISPers CA21106 and BridgeQG CA23130, supported by COST (European Cooperation in Science and Technology). The work of A.D. is supported by the Department of Atomic Energy, Government of India, under Project Identification Number RTI 4002.
Appendix A Long-range scalar field outside a compact star
Here, we follow Goldreich:1969sb ; Shapiro:1983du to calculate the electric and magnetic field profiles for a compact star when its spin axis aligns with its magnetic dipole axis, to calculate analytic expressions that are valid even for fast-rotating stars, i.e., when $\Omega R\sim{\cal O}(1)$. The dipolar magnetic field outside of a compact star is given in Eq. 1. The magnetic field just inside the surface of the star is given as
[…]
If $\bm{\mathcal{J}}$ denotes the current density then Ohm’s law reads $\bm{\mathcal{J}}=\sigma(\mathbf{E}^{\mathrm{in}}+\mathbf{v}\times\mathbf{B}^{\mathrm{in}})$, where $\sigma$ denotes the conductivity and $\mathbf{v}$ denotes the velocity of the star. Assuming that the NS is a perfect conductor ($\bm{\mathcal{J}}/\sigma\rightarrow 0$), we can write the Ohm’s law as $\mathbf{E}^{\mathrm{in}}+(\mathbf{\Omega}\times\mathbf{r})\times\mathbf{B}^{\mathrm{in}}=0$, since $\mathbf{v}=\mathbf{\Omega}\times\mathbf{r}$. Using Eq. 39, we obtain the electric field just inside the surface of the NS as
$\mathbf{E}^{\mathrm{in}}_{(r=R)}=-B_{0}\Big[\cos\theta\,(\mathbf{\Omega}\times\mathbf{r})\times\hat{r}+\frac{\sin\theta}{2}(\mathbf{\Omega}\times\mathbf{r})\times\hat{\theta}\Big$]. (40)
For $\mathbf{v}=\Omega R\sin\theta\,\hat{\phi}$, Eq. 40 becomes
[…]
Since the tangential component of the electric field is continuous at $r=R$, from Eq. 41 we obtain
$\mathbf{E}^{\mathrm{out}}_{\theta(r=R)}=-\frac{\partial}{\partial\theta}\Big(\frac{B_{0}\Omega R\sin^{2}\theta}{2}\Big)=\frac{\partial}{\partial\theta}\Big(\frac{B_{0}\Omega R}{3}P_{2}(\cos\theta)\Big),$ (42)
where $P_{2}(\cos\theta)=\frac{1}{2}(3\cos^{2}\theta-1)$ is the Legendre polynomial of degree $2$. Assuming the outer region of the star is vacuum, we can write $\mathbf{E}^{\mathrm{out}}=-\nabla\Phi$, where $\nabla^{2}\Phi=0$ from Poisson’s equation. Using the boundary condition Eq. 42 at $r=R$, the solution of Poisson’s equation becomes
$\Phi=-\frac{B_{0}\Omega R^{5}}{3r^{3}}P_{2}(\cos\theta).$ (43)
Thus, the scalar potential is quadrupolar in nature. Using Eq. 43, we obtain the expression for the electric field profile outside of the compact star as given in Eq. 2.
To solve Eq. 5 when $\Omega$ is not negligibly small, it is crucial to include the contribution from $\mathbf{E}^{2}$, since $\mathbf{E}^{2}$ in Eq. 3 contains terms proportional to $\Omega^{2}$. Therefore, we use the Green’s function method in solving the inhomogeneous differential equation Eq. 5. The source term is given as
$J(r,\theta)=-g_{\phi\gamma\gamma}\frac{B^{2}_{0}R^{6}}{4r^{6}}(3\cos^{2}\theta+1)+g_{\phi\gamma\gamma}\frac{B^{2}_{0}\Omega^{2}R^{10}}{4r^{8}}(5\cos^{4}\theta-2\cos^{2}\theta+1).$ (44)
The static Green’s function $G(x,y)$ satisfies
$\nabla^{2}G(x,y)=-\delta^{3}(x-y)/\sqrt{g(y)},$ (45)
and one can obtain the solution of the scalar field as
$\phi(x)=-\int d^{3}y\sqrt{g(y)}G(x,y)J(y).$ (46)
We can write Eq. 45 for a point source in a Schwarzschild background at $r=b$ and $\theta_{0}=\varphi_{0}=0$, as
$\begin{split}\frac{1}{r^{2}}\frac{\partial}{\partial r}\Big[(r^{2}-2Mr)\frac{\partial G}{\partial r}\Big]+\frac{1}{r^{2}\sin\theta}\frac{\partial}{\partial\theta}\Big[\sin\theta\frac{\partial G}{\partial\theta}\Big]=-\frac{\delta(r-b)\delta(\cos\theta_{0}-1)\delta(\varphi_{0})}{r^{2}}.\end{split}$ (47)
The solution of this homogeneous equation in terms of spherical harmonics can be written as
$G(r,\theta)=\sum_{l}R_{l}(r)P_{l}(\cos\theta),$ (48)
where
$\frac{\partial}{\partial r}\Big[(r^{2}-2Mr)\frac{\partial R_{l}}{\partial r}\Big]-l(l+1)R_{l}=0.$ (49)
Therefore, considering that the scalar field is finite at $r\rightarrow\infty$ and at $r\rightarrow 2M$, and continuous at $r=b$, we obtain the solution of the Green’s function as Campbell:1991kz
$G(r,\theta)$ $=$ $\sum^{\infty}_{l=0}C_{l}P_{l}\Big(\frac{b-M}{M}\Big)Q_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta),~~~r>b\ ,$ (50)
$=$ $\sum^{\infty}_{l=0}C_{l}Q_{l}\Big(\frac{b-M}{M}\Big)P_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta),~~~r<b\ ,$
where $P_{l}$ and $Q_{l}$ denote the Legendre polynomials of degree $l$ of first and second kind, respectively, and $C_{l}=(2l+1)/(4\pi M)$.
Hence, the external scalar field solution in terms of the Green’s function becomes
$\phi(r,\theta)=-\int^{\infty}_{r_{s}}db\int^{\pi}_{0}d\theta_{0}\int^{2\pi}_{0}d\varphi_{0}b^{2}\sin\theta_{0}G(r,\theta,\varphi,b,\theta_{0},\varphi_{0})J(b,\theta_{0},\varphi_{0}),$ (51)
where $r_{s}=2M$. Since the source term does not depend on $\varphi_{0}$, we can immediately perform the integration for $\varphi_{0}$ and write Eq. 51 as
$\begin{split}\phi(r,\theta)=-\sum_{l=0}^{\infty}\frac{2l+1}{2M}\int^{r}_{2M}db\int^{\pi}_{0}d\theta_{0}b^{2}\sin\theta_{0}P_{l}\Big(\frac{b-M}{M}\Big)Q_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta)P_{l}(\cos\theta_{0})J(b,\theta_{0})\\ -\sum^{\infty}_{l=0}\frac{2l+1}{2M}\int^{\infty}_{r}db\int^{\pi}_{0}d\theta_{0}b^{2}\sin\theta_{0}Q_{l}\Big(\frac{b-M}{M}\Big)P_{l}\Big(\frac{r-M}{M}\Big)P_{l}(\cos\theta)P_{l}(\cos\theta_{0})J(b,\theta_{0}).\end{split}$ (52)
Evaluating the integrals in Eq. 52, we obtain the scalar field profile outside the rotating star $(r>R)$ as
$\phi(r)\approx-\frac{g_{\phi\gamma\gamma}B^{2}_{0}\Omega^{2}R^{10}}{480M^{5}r}+\frac{g_{\phi\gamma\gamma}B_{0}^{2}R^{6}}{48M^{3}r}+\mathcal{O}\Big(\frac{1}{r^{2}}\Big).$ (53)
The dominant term of the scalar field is the monopole term $(l=0)$ and we can write the scalar field configuration as $\phi(r)\approx Q_{\phi}^{\mathrm{K}}/r$, where $Q_{\phi}^{\mathrm{K}}$ is the scalar charge, defined as
$Q_{\phi}^{\mathrm{K}}=-\frac{g_{\phi\gamma\gamma}B^{2}_{0}\Omega^{2}R^{10}}{480M^{5}}+\frac{g_{\phi\gamma\gamma}B^{2}_{0}R^{6}}{48M^{3}}.$ (54)
In the limit $\Omega R\ll 1$, Eqs. 53 and 54 reduce to Eqs. 7 and 8 respectively.
The scalar-induced magnetic field in the large $\Omega$ limit can similarly be obtained by solving Eq. 13 as
[…]
where $Q_{\phi}^{\mathrm{K}}$ is given in Eq. 54. The limiting scenario where $\Omega R\ll 1$ from Eq. 55 is given by Eq. 14. Therefore, the scalar-induced magnetic field is dipolar $(l=1)$ along the radial direction and monopolar $(l=0)$ along the angular direction.